CS 511: Organization of Programming Languages

Winter Quarter (15-625-511-001) Meets: M-W-F 8:00–8:50 in Swift 605

Instructor: Fred Annexstein Office: 889 Rhodes

Office Hours: M-W-F 9:00–10:00 Telephone: 556-1807

Description: The student shall learn the fundamental tools needed for critical evaluation of existing and future programming languages and constructs. The material taught in this course also is helpful when learning compiler design and construction. It is assumed that the student has a good working knowledge of the “C” language.

Prerequisites: See Course Descriptions in Department Office.

Grading (approx) : Homeworks count 50%, the midterm counts 25% and the final 25%. Class participation helps to bias things a bit. Grades are assigned on an informal “curve”.

Homework Policy: Six to seven homeworks may be assigned this quarter.

Textbook: Programming Languages: Concepts and Constructs by R. Sethi, Addison-Wesley (1996)

Accounts: Accounts will be on UCUNIX. Your account name and initial password are derived as usual.

Joint Work: Students may discuss homework assignments with each other, but must write solutions individually.

Approximate Schedule, SUBJECT TO CHANGE

Week  Class Material  Reading
1  Introduction  Chapter 1
2  Basic syntax, lexical scoping Chapter 2
3  Structured Programming Chapter 3
4  Data Representation  Chapter 4
5  Procedures  Chapters 5,6,7
6  Review, Midterm
7  Functional  Chapter 8
8  Use of Types, Lists  Chapter 9,10
9  Logic Programming  Chapter 11
10  Other Topics